# Towards Privacy-aware Mental Health AI Models: Advances, Challenges, and Opportunities

### v2

@@ Conclusion
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. This work has also been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). This research was further supported by the DYNAMIC center, which is funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number: LOEWE/1/16/519/03/09.001(0009)/98). T.C. acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 951847 and Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence. T.C also acknowledges the support of the Alexander von Humboldt Foundation through a Humboldt Research Fellowship for Experienced Researchers.

### v3

@@ Conclusion
This research work of I.G. and A.M. has been supported by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE, by the LOEWE Chair of Research Excellence “Ubiquitous Knowledge Processing” (Grant Number: LOEWE/4a//519/05/00.002(0002)/81)and by the DYNAMIC center, both of the former funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number: LOEWE/1/16/519/03/09.001(0009)/98). T.C. acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 951847 and Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence and the support of the Alexander von Humboldt Foundation through a Humboldt Research Fellowship for Experienced Researchers.